Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis of standard direct labor-hours. The budgeted variable manufacturing overhead is $3.20 per direct labor-hour and the budgeted fixed manufacturing overhead is $864,000 per year. 


The standard quantity of materials is 4 pounds per unit and the standard cost is $6.00 per pound. The standard direct labor-hours per unit is 1.5 hours and the standard labor rate is $12.60 per hour. 


The company planned to operate at a denominator activity level of 120,000 direct labor-hours and to produce 80,000 units of product during the most recent year. Actual activity and costs for the year were as follows: 

Actual number of units produced96,000
Actual direct labor-hours worked156,000
Actual variable manufacturing overhead cost incurred$312,000
Actual fixed manufacturing overhead cost incurred$936,000


Required:

 1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements.

 2. Prepare a standard cost card for the company's product.

 3a. Compute the standard direct labor-hours allowed for the year's production.

 3b. Complete the following Manufacturing Overhead T-account for the year.

 4. Determine the reason for any underapplied or overapplied overhead for the year by computing the variable overhead rate and efficiency variances and the fixed overhead budget and volume variances.
Answer #1

Solution 1:

Predetermined overhead rate for the year = Budgeted overhead / Budgeted labor hours

= [(120000*$3.20) + $864,000) / 120000 = $10.40 per direct labor hour

Variable overhead rate per labor hour = $3.20

Fixed overhead rate per labor hour = $7.20

Solution 2:

Standard cost Card - Lane Company
Particulars Qty Rate Per unit
Direct material 4 Pound at $6.00 Per Pound $24.00
Direct labor 1.5 DLHs at $12.60 Per DLH $18.90
Variable overhead 1.5 DLHs at $3.20 Per DLH $4.80
Fixed overhead 1.5 DLHs at $7.20 Per DLH $10.80
Standard cost per unit $58.50

solution 3a:

Standard direct labor-hours allowed for the year’s production = 96000*1.50 = 144000 hours

Solution 3b:

Manufacturing overhead - Lane Company
Particulars Debit Particulars Credit
Variable overhead incurred $312,000.00 Applied overhead (96000*1.5*$10.40) $1,497,600.00
Fixed overhead incurred $936,000.00
Overapplied overhead $249,600.00
Total $1,497,600.00 Total $1,497,600.00

Solution 4:

Actual rate of variable overhead = $312,000 / 156000 = $2 per labor hour

Variable overhead rate variance = (SR - AR) * AH = ($3.20 - $2) * 156000 = $187,200 F

Variable overhead efficiency variance = (SH - AH) *SR = (144000 - 156000) * $3.20 = $38,400 U

Fixed overhead budget variance = Budgeted fixed overhead - Actual fixed overhead = $864,000 - $936,000 = $72,000 U

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead = (144000*$7.20) - $864,000

= $172,800 F